OpenStudy (anonymous):

A particle's position as a function of time t is given by r⃗ =(5.0t+6.0t2)mi+(7.0−3.0t3)mj. At t=5.0s, find the magnitude of the particle's displacement vector Δr⃗ relative to the point r⃗ 0=(0.0i+7.0j)m. At t=5.0s, find the angle of the particle's displacement vector Δr⃗ relative to the point r⃗ 0=(0.0i+7.0j)m. Help please i'm stuck

OpenStudy (loser66):

\(\vec r_0\) means the position of the particle at time t =0 . It is \(\vec r_0= (0.0 i + 7.0 j)m\) at t=5.0s, the position of it is \(\vec r_5= (5.0*5) i + (6.0*(5^2))j= (25.0 i +150 j)\) --------------------------------------- \(\triangle r =r_5 - r_0 = 25.0i +143 j)\)